Grade 11Trigonometry

Angle of triangle are in ap and the number of degrees in least angle is to the number of degrees in mean angle is 2:3 find angle in radian

Deepika dubey
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Let angles are 
a-d, a, a+d
a-d +a + a+d = 180
3a = 180
a= 60°
Now 
a-d/a = 2/3
3a - 3d = 2a
a = 3d
Hence d = 20
Hence angles are 40° , 60° and 80°
Now you can convert them in radians